# How Well do LLMs Compress Their Own Chain-of-Thought? A Token Complexity Approach

### v1

@@ Abstract
Chain-of-thought prompting has emerged as a powerful technique for enabling large language models (LLMs) to solve complex reasoning tasks. However, these reasoning chains can be verbose, raising concerns about efficiency. In response, recent works have sought to decrease response lengths through simple prompting strategies (e.g. ‘be concise’). In this work, we conduct the first systematic study of the relationship between reasoning length and model performance across a diverse range of compression instructions (e.g. ‘use 10 words or less’ or ’remove all punctuation’). In doing so, we discover a universal tradeoff between reasoning length and accuracy that persists across even very distinct reasoning chains. We demonstrate that this tradeoff emerges from a sharp threshold behavior at the question level: each task has an intrinsic ‘token complexity’ – a minimal number of tokens required for successful problem-solving. We show how token complexity enables us to compute information-theoretic limits on the accuracy-compression tradeoff, and find that prompt-based compression strategies operate far from these theoretical limits. This suggests there may be significant room for improvement and our framework provides a benchmark to help researchers evaluate progress in reasoning efficiency. Our work also highlights the importance of adaptive compression – giving shorter responses for easier questions – and we show that token complexity is a useful tool for measuring this capability.111 Our code is available at https://github.com/Compressed-CoT/compressed-cot. Our data is available at https://huggingface.co/datasets/cot-compression/TokenComplexity
How Well do LLMs Compress Their Own Chain-of-Thought?
A Token Complexity Approach
Ayeong Lee Columbia Business School al3393@columbia.edu                        Ethan Che Columbia Business School ewc2119@columbia.edu                        Tianyi Peng Columbia Business School tp2845@columbia.edu
@@ 1 Introduction
Recent advancements in large language models (LLMs)—including GPT-o1 and DeepSeek R1—alongside broader AI agent development, have showcased impressive reasoning capabilities, hinting at a future where complex problem-solving and decision-making can be automated. However, this rapid progress also introduces a significant challenge: the computational cost of reasoning is projected to increase substantially as these models are deployed in real-world applications. This growing inference cost highlights the necessity for efficient reasoning strategies, motivating our research into reducing the computational burden of LLM inference while maintaining high performance.
A pivotal technique for enhancing LLM reasoning has been chain-of-thought (CoT) prompting, which encourages models to generate intermediate reasoning steps before arriving at a final answer Wei et al. (2022); Kojima et al. (2022). While effective, these reasoning chains often involve lengthy intermediate computations, significantly increasing inference costs when deployed at scale Yu et al. (2024).
Although CoT-generated outputs are typically verbose, containing redundant Chiang and Lee (2024) or sometimes irrelevant information Wang et al. (2023), it remains unclear how best to compress the chain-of-thought content for effective problem-solving. Prior work has observed that asking the LLM to ‘be concise’ or ‘use at most 100 words’ can reduce response length while incurring a range of degradation to accuracy Jin et al. (2024); Renze and Guven (2024); Han et al. (2024). Yet so far, there has not been a comprehensive exploration of the effectiveness of different compression strategies. For example, should language models use fewer reasoning steps instead of using fewer words or characters? Should they be removing unnecessary punctuation or filler words (e.g. ‘therefore’, ‘so’). Should they reason in alternative languages that may be more token-efficient?
@@
BeConcise (…), but be concise. Let’s solve this step by step: 1) Z18 (…)
BulletPoints (…) only use bullet points. - Z18 is the group of integers modulo 18 (…)
OnlyNumbers (…) only use numbers or equations. |Z18| = 18. Answer: (G)
NoSpaces (…) do not use any spaces or line breaks. TheorderofagroupZnisthesamenumbern,sothe (…)
@@
NoProperGrammar (…) do not use proper grammar. lemme help u with this z18 is group
AbbreviateWords (…) abbreviate words as much as possible. Slvng fr rdr f Z18: Z18 = grp f ntgrs mdle (…)
WordLimit(k) (…) use at most $k$ words. $k\in\{1,...,100\}$ Order of Z18 is eighteen. Answer: (G)
CharLimit(k) (…) use at most $k$ letters. $k\in\{1,...,500\}$ The order of a group is the number of elements in (…)
TokenLimit(k) (…) use at most $k$ tokens. $k\in\{1,...,500\}$ The order of Z18 is 18, as it contains elements (…)
StepLimit(k) (…) use at most $k$ steps. $k\in\{1,...,5\}$ Step 1: The order of a group Zn is simply n, so (…)
ChineseCoT (…) Respond in Chinese {CJK} UTF8gbsn让我帮你解答这个问题. Z18 是-个循环群 (…)
ChineseCoT(k) (…) Use at most $k$ Chinese characters. $k\in\{1,...,500\}$CJK} UTF8gbsn答案是18. Answer: (G).
@@
Table 1: The Chain-of-Thought prompts we consider. The right column gives an example of the chain-of-thought of Claude 3.5 Sonnet on a sample problem in MMLU-Pro Math.
@@
The main contribution of our work is to provide the first systematic study of the trade-off between reasoning length and performance across different prompt-based compression strategies, including prior strategies such as ‘be concise’ as well as alternative approaches such as ‘only use bullet points’ or ‘use at most 50 Chinese characters’. In total, we evaluate 31 prompts for six LLMs on three standard reasoning datasets. Remarkably, although these prompting strategies produce different chains of thought, their trade-offs between response length and accuracy lie on a universal trade-off curve. In other words, all prompts are equally "good" as extremes on this curve. What primarily affects accuracy is the length of the chain of thought, far more than changes in its composition.
@@
Our second contribution is a novel empirical observation: the performance of reasoning tasks exhibits a sharp threshold dependence on reasoning length at the question level. By evaluating multiple prompts for each question, we demonstrate that most questions have a well-defined ‘token complexity’—a minimum number of tokens required to successfully solve the question. This token complexity is universal across prompting strategies. We estimate token complexities across various benchmarks and find that: (i) Token complexity alone can predict the performance of CoT prompting strategies with 95% accuracy. (ii) It serves as a robust measure of reasoning task difficulty, enabling us to investigate whether LLMs reason adaptively—using shorter chains-of-thought for easier questions.
These results raise the question of whether the trade-off curve between response-length and accuracy induced by these prompting strategies are close or far from optimal. Viewing these strategies as a form of ‘lossy compression’, we borrow tools from rate-distortion theory to characterize the optimal accuracy-compression trade-off. In doing so, we find that prompt-based strategies are far from the theoretical limits, especially on harder datasets. In particular, our analysis highlights the importance of adaptive compression and shows that while LLMs tend to use shorter reasoning chains for easier problems, there is significant room for improvement. In total, this work provides a set of empirical tools that can help researchers benchmark new methodologies for improving the efficiency of LLM reasoning.
@@ 1.1 Related Work
Recent research has begun to gain traction in exploring the trade-off between response length and accuracy in LLMs. Studies such as Renze and Guven (2024); Jin et al. (2024); Nayab et al. (2024); Han et al. (2024) have employed specific prompting strategies to constrain response length and assess the associated impact on accuracy and performance. Additionally, several works have highlighted the redundancy inherent in CoT prompting Chiang and Lee (2024); Wu et al. (2025) and emphasized the benefits of concise reasoning. Other approaches have focused on fine-tuning strategies to adapt LLMs for generating more succinct reasoning Kang et al. (2024); Yu et al. (2024).
Our work advances this growing body of literature by making three key contributions: 1) We conducted a systematic evaluation of a rich set of prompts designed to reduce the length of CoT reasoning while maintaining accuracy; 2) We find that accurate answers are only achieved when the output length exceeds a certain threshold, which is intrinsic to the problem and independent of the CoT format. We formalize this concept as the token complexity of a problem; 3) We derive theoretical limits on the length-accuracy tradeoff, providing a framework for researchers to benchmark new methodologies aimed at compressing chain-of-thought reasoning effectively.
@@ 2 Experiments
Our evaluation encompasses the following LLMs: GPT-4o Hurst et al. (2024), GPT-4o-mini Hurst et al. (2024), Claude 3.5 Sonnet Anthropic (2024), Claude 3.5 Haiku Anthropic (2024), and Llama 3.3 70B Instruct Dubey et al. (2024). We evaluate these models on three standard math reasoning datasets: MATH-500 Lightman et al. (2023), a random 500 problem subset of GSM8K (GSM8K, Cobbe et al. (2021)), and a random 500 problem subset of MMLU-Pro Math problems (MMLU-Pro Math, Wang et al. (2024)).
For each LLM and dataset, we test 31 prompts designed to induce shorter response lengths, detailed in Table 1. These prompts include ones considered in prior literature: ‘be concise’ Renze and Guven (2024), ‘use $k$ words or less’ Jin et al. (2024); Nayab et al. (2024), ‘use $k$ tokens or less’ Han et al. (2024), but include additional curated ones to assess the impact of alternative compression strategies. For each prompt, we assess performance with two metrics: (1) accuracy, the fraction of questions solved correctly, and (2) average token length, the average number of output tokens produced by the LLM in their response across questions in the dataset.
Model Prompt Accuracy Token Count GPT-4o NoCoT 42% 10 DefaultCoT 81% 586 BeConcise 80% 415 BulletPoints 75% 185 OnlyNumbers 79% 222 NoSpaces 79% 249 NoProperGrammar 79% 189 AbbreviateWords 74% 276 ChineseCoT 75% 326 Claude 3.5 Sonnet NoCoT 48% 8 DefaultCoT 81% 324 BeConcise 78% 227 BulletPoints 76% 167 OnlyNumbers 77% 130 NoSpaces 77% 188 NoProperGrammar 77% 173 AbbreviateWords 77% 220 ChineseCoT 76% 311 Llama-3.3-70B NoCoT 43% 20 DefaultCoT 74% 552 BeConcise 77% 443 BulletPoints 70% 194 OnlyNumbers 63% 304 NoSpaces 67% 330 NoProperGrammar 68% 217 AbbreviateWords 70% 327 ChineseCoT 74% 363
Figure 1: For each of the 31 CoT prompts we consider (see legend above), we report average token length vs accuracy for GPT-4o and Claude 3.5 Sonnet on MMLU-Pro Math and GPT-4o-mini on GSM8K tasks. Despite differences in the chain-of-thought, many of them live on a universal tradeoff curve. We report upper bounds for optimal accuracy under a given average token budget (see section 4). See Appendix E for more models and benchmarks.
@@ 2.1 Benchmark Results
By considering multiple diverse prompts, we are able to induce chains-of-thought along a spectrum of response lengths and reasoning performance, with NoCoT (no chain-of-thought) using the fewest tokens with the lowest accuracy and DefaultCoT (i.e. standard chain-of-though-prompting ‘think step-by-step’) using the most tokens and generally having the highest benchmark performance. The dataset is available at https://huggingface.co/datasets/cot-compression/TokenComplexity. In Table 2, we focus on a subset of chain-of-thought prompts and report the accuracy and average token count for MMLU-Pro Math across several LLMs. We highlight the following observations, which also hold for other datasets (see Appendix A).
@@
  • BeConcise Renze and Guven (2024) consistently reduces token length without significantly hurting performance.
  • Yet, there are other prompts such as OnlyNumbers or NoProperGrammar that preserve a similar accuracy as BeConcise but induce substantially shorter chains-of-thought (a $\approx 50\%$ reduction for GPT-4o). The strong performance of these prompts suggests that LLMs do not necessarily require proper English to conduct chain-of-thought reasoning effectively.
  • There is no universally dominant prompt; while OnlyNumbers and NoProperGrammar generally work well for GPT-4o and Claude 3.5 Sonnet, they are less performant for LLaMA 3.3 70B.
  • While LLMs do not always follow the prompt exactly, the prompt still induces large reductions in the response length (as also observed in  Han et al. (2024)).
@@ 2.2 Universal Trade-off between Reasoning Length and Accuracy
Figure 2: Evidence of the token complexity hypothesis. (Left) Performance of Claude 3.5 Sonnet on a sample question in MATH-500 exhibits a threshold behavior. With the exception of 2 prompts, all the prompts that use more tokens than the threshold get the answer correct, the rest do not. The red dotted line indicates the estimated token complexity $\tau_{i}^{\pi}$ from the results. (Right) Actual accuracy on MATH-500 vs predicted accuracy from the token complexity hypothesis. Token complexity is highly predictive of actual accuracy.
@@
To study this question, we plot the average token-length and accuracy of all 31 prompts in Figure 1 for the MMLU-Pro Math benchmark (results for other benchmarks are in Appendix E). Remarkably, we see that almost all the prompts we consider lie on a universal trade-off curve between response length and accuracy. This suggests that regardless of whether the chain-of-thought is formatted in bullet points, without spaces, using only numbers, or even in Chinese, ultimately it is the length of the chain-of-thought that matters most. This result also holds for the Wordlimit(k), Charlimit(k), TokenLimit(k), StepLimit(k), and ChineseCoT(k) prompts, which ask the LLM to limit the response to be at most $k$ words, letters, tokens, reasoning steps, or Chinese characters. One would expect that each of these strategies would result in separate tradeoff curves, but surprisingly all of these strategies result in a near-identical trade-off between response length and accuracy. This suggests that there is not much room for improvement simply by changing the composition of the chains-of-thought.
@@ 3 The Token Complexity Hypothesis
As discussed before, we should think of this hypothesis as holding only for ‘reasonable’ chains-of-thought (e.g. not 100 tokens of pure whitespace). Nonetheless, as we will discuss, we observe this to hold broadly for all the prompts we test. This universality makes the token complexity notion useful because it provides fine-grained notion of problem difficulty that doesn’t depend on a specific prompting strategy: more difficult problems simply require more tokens to solve.
@@
We proceed to test this hypothesis quantitatively across LLMs $\pi$ and benchmarks. First, we measure to what degree success or failure on a task can be classified based purely on the token-length of the chain-of-thought, i.e. whether the behavior in the left panel of Figure 2 holds broadly. To do so, we use our dataset of $k=1,...,K=31$ chain-of-thought prompts for each question. For $t\in\mathbb{N}$, we define $c^{\pi}_{i}(t)$ to be the classification accuracy under a threshold classifier under $t$:
$\hat{\tau}^{\pi}_{i}$ $\equiv\arg\max_{k}c_{i}^{\pi}(t(X_{i,k}))$ (1)
$c_{i}^{*}$ $\equiv\max_{k}c_{i}^{\pi}(t(X_{i,k}))$ (2)
$\bar{c}_{\pi}$ $\equiv\frac{1}{n}\sum\nolimits_{i=1}^{n}c_{i}^{*}$ (3)
We let $\hat{\tau}_{i}^{\pi}=\infty$ if setting the threshold to $t=\infty$ results in better classification accuracy, e.g. if none of the chains-of-thoughts correctly solve the problem in which case $a_{i}^{\pi}(X_{i,k})=0$ for all $k$. For each dataset and model, we report the average classification accuracy $\bar{c}_{\pi}$ under the estimated $\tau^{\pi}_{i}$ in Table 4. Overall, the average classification accuracy is very high, above 90% for most models and benchmarks, verifying that (1) token-length is highly predictive for performance at the question-level and (2) question-level performance exhibits a threshold relationship with token-length. And we see that the threshold classifier achieves higher average classification accuracy for (1) more capable models (i.e. GPT-4o and Claude 3.5 Sonnet) and (2) easier benchmarks (i.e. GSM8K), which mirrors the finding in Section 2.2 concerning adherence to the universal trade-off curve.
$\text{Acc}_{\pi}(P_{k})$ $\equiv\frac{1}{n}\sum\nolimits_{i=1}^{n}a^{\pi}_{i}(X_{i,k})$ (4)
$\widehat{\text{Acc}}_{\pi}(P_{k})$ $\equiv\frac{1}{n}\sum\nolimits_{i=1}^{n}\mathbf{1}\{t(X_{i,k})\geq\hat{\tau}^{\pi}_{i}\}$ (5)
$\text{Err}_{\pi}=\frac{1}{K}$ $\sum\nolimits_{k=1}^{K}\frac{|\text{Acc}_{\pi}(P_{k})-\widehat{\text{Acc}}_{\pi}(P_{k})|}{\text{Acc}_{\pi}(P_{k})}$ (6)
In Table 5, for each LLM and dataset we report $\text{Err}_{\pi}$, the average relative discrepancy between actual accuracy and predicted accuracy. We observe that the token threshold classifier is able to predict overall benchmark performance within 6% error, and the error is even smaller for larger models on easier benchmarks. This illustrates that the token-complexity hypothesis provides a model for reasoning performance that is highly predictive of actual performance, which can be seen visually in the right panel of Figure 2 for Claude 3.5 Sonnet on the MATH-500 dataset. Since predicted accuracy $\widehat{\text{Acc}}_{\pi}(P_{k})$ only depends on the token lengths, this also provides an explanation for why accuracy across the chain-of-thought prompts we consider is highly dependent on average token-length in Figure 1. We explore the implications of this hypothesis in the next section.

### v2

@@ Abstract
Chain-of-thought prompting has emerged as a powerful technique for enabling large language models (LLMs) to solve complex reasoning tasks. However, these reasoning chains can be verbose, raising concerns about efficiency. In response, recent works have sought to decrease response lengths through simple prompting strategies (e.g. ‘be concise’). In this work, we conduct the first systematic study of the relationship between reasoning length and model performance across a diverse range of compression instructions (e.g. ‘use 10 words or less’ or ’remove all punctuation’). In doing so, we discover a universal tradeoff between reasoning length and accuracy that persists across even very distinct reasoning chains. We demonstrate that this tradeoff emerges from a sharp threshold behavior at the question level: each task has an intrinsic ‘token complexity’ – a minimal number of tokens required for successful problem-solving. We show how token complexity enables us to compute upper bounds on the optimal accuracy-compression tradeoff. Our analysis reveals that prompt-based compression strategies operate far from these theoretical limits, suggesting significant room for improvement and providing benchmarks to help researchers evaluate progress in reasoning efficiency. Our work also highlights the importance of adaptive compression – giving shorter responses for easier questions – and we show that token complexity is a useful tool for measuring this capability.
@@ 1 Introduction
Recent advancements in large language models (LLMs)—including o1 and DeepSeek R1—alongside broader AI agent development, have showcased impressive reasoning capabilities, hinting at a future where complex problem-solving and decision-making can be automated. However, this rapid progress also introduces a significant challenge: the computational cost of reasoning is projected to increase substantially as these models are deployed in real-world applications. This growing inference cost highlights the necessity for efficient reasoning strategies, motivating our research into reducing the computational burden of LLM inference while maintaining high performance.
A pivotal technique for enhancing LLM reasoning has been chain-of-thought (CoT) prompting, which encourages models to generate intermediate reasoning steps before arriving at a final answer (Wei et al., 2022; Kojima et al., 2022). While effective, these reasoning chains often involve lengthy intermediate computations, increasing inference costs when deployed at scale (Yu et al., 2024).
Although CoT-generated outputs are typically verbose, containing redundant (Chiang & Lee, 2024) or sometimes irrelevant information (Wang et al., 2023), it remains unclear how best to compress the chain-of-thought content for effective problem-solving. Prior work has observed that asking the LLM to ‘be concise’ or ‘use at most 100 words’ can reduce response length while incurring a range of degradation to accuracy (Jin et al., 2024; Renze & Guven, 2024; Han et al., 2024). Yet so far, there has not been a comprehensive exploration of the effectiveness of different compression strategies. For example, should language models use fewer reasoning steps instead of using fewer words or characters? Should they be removing unnecessary punctuation or filler words (e.g. ‘therefore’, ‘so’). Should they reason in alternative languages that may be more token-efficient?
@@
BeConcise Be concise. Let’s solve this step by step: 1) Z18 (…)
BulletPoints Only use bullet points. - Z18 is the group of integers modulo (…)
OnlyNumbers Only use numbers or equations. —Z18— = 18. Answer: (G)
NoSpaces Do not use any spaces or line breaks. TheorderofagroupZnisthesame (…)
@@
NoProperGrammar Do not use proper grammar. lemme help u with this z18 is group
AbbreviateWords Abbreviate words as much as possible. Slvng fr rdr f Z18: Z18 = grp f ntgrs (…)
WordLimit(k) Use at most $k$ words. $(k\in[1,100])$ Order of Z18 is eighteen. Answer: (G)
CharLimit(k) Use at most $k$ letters. $(k\in[1,500])$ The order of a group is the number of (…)
TokenLimit(k) Use at most $k$ tokens. $(k\in[1,500])$ The order of Z18 is 18, as it contains (…)
StepLimit(k) Use at most $k$ steps. $(k\in[1,5])$ Step 1: The order of a group Zn is (…)
ChineseCoT Respond in Chinese {CJK} UTF8gbsn让我帮你解答这个问题 (…)
ChineseCoT(k) Use at most $k$ Chinese characters. {CJK} UTF8gbsn答案是18. Answer: (G).
@@
Table 1: The Chain-of-Thought prompts we consider. The right column gives an example of the chain-of-thought of Claude 3.5 Sonnet on a sample problem in MMLU-Pro Math. We indicate the range of $k$ that we consider.
@@
The main contribution of our work is to provide the first systematic study of the trade-off between reasoning length and performance across different prompt-based compression strategies, including prior strategies such as ‘be concise’ as well as alternative approaches such as ‘only use bullet points’ or ‘use at most 50 Chinese characters’. In total, we evaluate 31 prompts for six LLMs on three standard reasoning datasets. Remarkably, although these prompting strategies produce different chains of thought, their trade-offs between response length and accuracy lie on a universal trade-off curve. In other words, all prompts are equally ”good” as extremes on this curve. What primarily affects accuracy is the length of the chain of thought, far more than changes in its composition.
@@
Our second contribution is a novel empirical observation: the performance of reasoning tasks exhibits a sharp threshold dependence on reasoning length at the question level. By evaluating multiple prompts for each question, we demonstrate that most questions have a well-defined ‘token complexity’—a minimum number of tokens required to successfully solve the question–which holds across diverse prompting strategies. We estimate token complexities across various benchmarks and find that: (i) Token complexity alone can predict the performance of CoT prompting strategies with 94% accuracy. (ii) It serves as a robust measure of reasoning task difficulty, enabling us to investigate whether LLMs reason adaptively—using shorter chains-of-thought for easier questions.
These results raise the question of whether the trade-off curve between response-length and accuracy induced by these prompting strategies are close or far from optimal. Viewing these strategies as a form of ‘lossy compression’, we take inspiration from rate-distortion theory to characterize an upper bound on the optimal accuracy-compression trade-off. In doing so, we find that prompt-based strategies are far from this upper bound, especially on harder datasets. We also illustrate how our experimental results can be used to evaluate methodologies for improving the adaptivity of LLM reasoning. We show evidence that even without explicitly prompting, LLMs do exhibit adaptive reasoning lengths and that the simple prompting strategies we test offer a surprisingly strong baseline that can outperform more sophisticated prompt-routing strategies. Finally, we show that if one has access to a verifier, one can achieve a substantially better accuracy-length tradeoff and approach the upper bound.
@@ 1.1 Related Work
Recent research has begun to gain traction in exploring the trade-off between response length and accuracy in LLMs. Studies such as (Renze & Guven, 2024; Jin et al., 2024; Nayab et al., 2024; Han et al., 2024) have employed specific prompting strategies to constrain response length and assess the associated impact on accuracy and performance. Additionally, several works have highlighted the redundancy inherent in CoT prompting (Chiang & Lee, 2024; Wu et al., 2025) and emphasized the benefits of concise reasoning. Other approaches have focused on fine-tuning strategies to adapt LLMs for generating more succinct reasoning (Kang et al., 2024; Yu et al., 2024).
Our work advances this growing body of literature by making three key contributions: (1) we conducted a systematic evaluation of a rich set of prompts designed to reduce the length of CoT reasoning while maintaining accuracy. (2) We find that accurate answers are only achieved when the output length exceeds a certain threshold, which is intrinsic to the problem and independent of the CoT format. We formalize this concept as the token complexity of a problem. (3) We derive theoretical limits on the length-accuracy tradeoff, providing a framework for researchers to benchmark new methodologies aimed at compressing chain-of-thought reasoning effectively.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 1: For each of the 31 CoT prompts we consider (see legend above), we report average token length vs accuracy for GPT-4o and Claude 3.5 Sonnet on MMLU-Pro Math and GPT-4o-mini on GSM8K tasks. Despite differences in the chain-of-thought, many of them live on a universal tradeoff curve. Using our framework, we compute upper bounds on optimal accuracy under a given average token budget (see section 4). See Appendix G for more models and benchmarks.
@@ 2 Experiments
Our evaluation encompasses the following LLMs: GPT-4o (Hurst et al., 2024), GPT-4o-mini (Hurst et al., 2024), Claude 3.5 Sonnet (Anthropic, 2024), Claude 3.5 Haiku (Anthropic, 2024), and Llama 3.3 70B Instruct (Dubey et al., 2024). We evaluate these models on three standard math reasoning datasets: MATH-500 (Lightman et al., 2023), a random 500 problem subset of GSM8K (GSM8K, Cobbe et al. (2021)), and a random 500 problem subset of MMLU-Pro Math problems (MMLU-Pro Math, Wang et al. (2024)).
For each LLM and dataset, we test 31 prompts designed to induce shorter response lengths, detailed in Table 1. These prompts include ones considered in prior literature: ‘be concise’ (Renze & Guven, 2024), ‘use $k$ words or less’ (Jin et al., 2024; Nayab et al., 2024), ‘use $k$ tokens or less’ (Han et al., 2024), but include additional curated ones to assess the impact of alternative compression strategies. For each prompt, we assess performance with two metrics: (1) accuracy, the fraction of questions solved correctly, and (2) average token length, the average number of output tokens produced by the LLM in their response across questions in the dataset.
@@ 2.1 Benchmark Results
By considering multiple diverse prompts, we are able to induce chains-of-thought along a spectrum of response lengths and reasoning performance, with NoCoT (no chain-of-thought) using the fewest tokens with the lowest accuracy and DefaultCoT (i.e. standard chain-of-though-prompting ‘think step-by-step’) using the most tokens and generally having the highest benchmark performance. In Table 2, we focus on a subset of chain-of-thought prompts and report the accuracy and average token count for MMLU-Pro Math across several LLMs. We highlight the following observations, which also hold for other datasets (see Appendix B).
@@
  • BeConcise (Renze & Guven, 2024) indeed consistently reduces token length without significantly hurting performance.
  • Yet, there are other prompts such as OnlyNumbers or NoProperGrammar that preserve a similar accuracy as BeConcise but induce shorter chains-of-thought (a $\approx 50\%$ reduction for GPT-4o). The strong performance of these prompts suggests that LLMs do not necessarily require proper English to conduct chain-of-thought reasoning effectively.
  • There is no universally dominant prompt; while OnlyNumbers and NoProperGrammar are effective for GPT-4o and Claude 3.5 Sonnet, they are less so for LLaMA 3.3 70B.
  • While LLMs do not always follow the prompt exactly, the prompt still induces large variations in the response length (as also observed in  Han et al. (2024)).
@@ 2.2 Universal Trade-off between Reasoning Length and Accuracy
Prompt GPT-4o Claude 3.5 Sonnet Llama-3.3-70B
Acc Tok Acc Tok Acc Tok
NoCoT 42% 10 48% 8 43% 20
DefaultCoT 81% 586 81% 324 74% 552
BeConcise 80% 415 78% 227 77% 443
BulletPoints 75% 185 76% 167 70% 194
OnlyNumbers 79% 222 77% 130 63% 304
NoProperGrammar 79% 189 77% 173 68% 217
ChineseCoT 75% 326 76% 311 74% 363
To study this question, we plot the average token-length and accuracy of all 31 prompts in Figure 1 for the MMLU-Pro Math benchmark (results for other benchmarks are in Appendix G). Remarkably, we see that almost all the prompts we consider lie on a universal trade-off curve between response length and accuracy. This suggests that regardless of whether the chain-of-thought is formatted in bullet points, without spaces, using only numbers, or even in Chinese, ultimately it is the length of the chain-of-thought that matters most. This result also holds for the Wordlimit(k), Charlimit(k), TokenLimit(k), StepLimit(k), and ChineseCoT(k) prompts, which ask the LLM to limit the response to be at most $k$ words, letters, tokens, reasoning steps, or Chinese characters. One would expect that each of these strategies would result in separate tradeoff curves, but surprisingly all of these strategies result in a near-identical trade-off between response length and accuracy. This suggests that there is not much room for improvement by changing the composition of the chains-of-thought alone.
@@ 3 The Token Complexity Hypothesis
Figure 2: Illustration of the token complexity hypothesis. (Left) Performance of Claude 3.5 Sonnet on a sample question in MATH-500 exhibits a threshold behavior. With the exception of 2 prompts, all the prompts that use more tokens than the threshold get the answer correct, the rest do not. The red dotted line indicates the estimated token complexity $\tau_{i}^{\pi}$ from the results. (Right) Actual benchmark accuracy on MATH-500 vs predicted accuracy from the token complexity hypothesis. Token complexity is highly predictive of actual accuracy (see Table 3).
@@
As discussed before, we should think of this hypothesis as holding only for ‘reasonable’ chains-of-thought (e.g. not 100 tokens of pure whitespace); and we observe this to hold broadly for all the prompts we test. This universality makes the token complexity notion useful because it provides fine-grained notion of problem difficulty that doesn’t depend on a specific prompting strategy: more difficult problems require more tokens to solve.
Model MMLU-Pro Math GSM8K MATH-500
$\bar{c}_{\pi}$ $\text{Err}_{\pi}$ $\bar{c}_{\pi}$ $\text{Err}_{\pi}$ $\bar{c}_{\pi}$ $\text{Err}_{\pi}$
GPT-4o 92.1% 5.0% 97.0% 1.6% 89.9% 6.0%
GPT-4o Mini 91.3% 5.2% 94.1% 5.1% 91.4% 5.7%
Claude 3.5 Sonnet 92.6% 3.4% 97.4% 1.4% 92.1% 4.5%
Claude 3.5 Haiku 90.1% 5.1% 94.8% 4.1% 90.2% 6.3%
Llama 3.3 70B 90.6% 5.1% 96.2% 2.7% 89.8% 6.1%
Table 3: Evidence of token complexity hypothesis across LLMs and reasoning benchmarks. $\bar{c}_{\pi}$ is the average classification accuracy of the threshold classifier $\mathbf{1}\{t(X_{i,k})>\hat{\tau}_{i}^{\pi}\}$ in predicting whether the LLM gets the question correct or not. $\text{Err}_{\pi}$ is the discrepancy between actual performance $\text{Acc}_{\pi}(P_{k})$ and predicted performance $\widehat{\text{Acc}}_{\pi}(P_{k})$ from the token complexity hypothesis. The high classification accuracy of $\bar{c}_{\pi}\approx 90\%$ and the small discrepancy $\text{Err}_{\pi}\approx 6\%$ illustrate that the token complexity hypothesis has strong predictive power for the reasoning performance of LLMs at both the question and benchmark levels.
We proceed to test this hypothesis quantitatively across LLMs $\pi$ and benchmarks. First, we measure to what degree success or failure on a task can be classified based purely on the token-length of the chain-of-thought, i.e. whether the behavior in the left panel of Figure 2 holds broadly. To do so, we use our dataset of $k=1,...,K=31$ chain-of-thought prompts for each question. For a threshold $t\in\mathbb{N}$, we define $c^{\pi}_{i}(t)$ to be the classification accuracy under a threshold classifier under $t$, which measures how predictable reasoning success is based on whether the token count exceeds threshold $t$
$\hat{\tau}^{\pi}_{i}$ $\equiv\arg\max_{k}c_{i}^{\pi}(t(X_{i,k})),\quad c_{i}^{*}\equiv\max_{k}c_{i}^{\pi}(t(X_{i,k})),\quad\bar{c}_{\pi}\equiv\frac{1}{n}\sum\nolimits_{i=1}^{n}c_{i}^{*}$ (1)
We let $\hat{\tau}_{i}^{\pi}=\infty$ if setting the threshold to $t=\infty$ results in better classification accuracy, e.g. if none of the chains-of-thoughts correctly solve the problem in which case $a_{i}^{\pi}(X_{i,k})=0$ for all $k$. For each dataset and model, we report the average classification accuracy $\bar{c}_{\pi}$ under the estimated $\tau^{\pi}_{i}$ in Table 4. Overall, the average classification accuracy is very high, above 90% for most models and benchmarks, verifying that (1) token-length is highly predictive for performance at the question-level and (2) question-level performance exhibits a threshold relationship with token-length. We also see that the threshold classifier achieves higher average classification accuracy for (1) more capable models (i.e. GPT-4o and Claude 3.5 Sonnet) and (2) easier benchmarks (i.e. GSM8K), which mirrors the finding in Section 2.2 concerning adherence to the universal trade-off curve.
$\text{Acc}_{\pi}(P_{k})\equiv\frac{1}{n}\sum\nolimits_{i=1}^{n}a^{\pi}_{i}(X_{i,k}),\quad\widehat{\text{Acc}}_{\pi}(P_{k})$ $\equiv\frac{1}{n}\sum\nolimits_{i=1}^{n}\mathbf{1}\{t(X_{i,k})\geq\hat{\tau}^{\pi}_{i}\}$ (2)
In Table 3, for each LLM and dataset we report $\text{Err}_{\pi}$, the average relative discrepancy between actual accuracy and predicted accuracy:
$\text{Err}_{\pi}=\frac{1}{K}\sum\nolimits_{k=1}^{K}\frac{|\text{Acc}_{\pi}(P_{k})-\widehat{\text{Acc}}_{\pi}(P_{k})|}{\text{Acc}_{\pi}(P_{k})}$
We observe that the token threshold classifier is able to predict overall benchmark performance within 6% error, and the error is even smaller for larger models on easier benchmarks. This illustrates that the token-complexity hypothesis provides an accurate model for reasoning performance, which can be seen visually in the right panel of Figure 2 for Claude 3.5 Sonnet on the MATH-500 dataset and helps characterize the strong dependence of token-length on accuracy. We explore the implications of this hypothesis in the next section.